OpenStudy (anonymous):

If B [7 3 -2] find -3b [5 4 3] [ 21 9 -6] [-15 -12 9] [21 9 -6] [15 12 9] [10 6 1] [8 7 6] [4 0 -5] [2 1 0 ]

OpenStudy (unklerhaukus):

multiply all elements of the matrix B, by -3

OpenStudy (anonymous):

can you give me a example ? @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\mathrm X = \begin{bmatrix}-2&3\\7&8\end{bmatrix}\] \[-5\mathrm X = -5\begin{bmatrix}-2&3\\7&8\end{bmatrix} =\begin{bmatrix}(-5\times-2)&(-5\times3)\\(-5\times7)&(-5\times8)\end{bmatrix} =\begin{bmatrix}10&-15\\-35&-40\end{bmatrix}\]

OpenStudy (unklerhaukus):

understand?

OpenStudy (anonymous):

Yes i see what you did there ? but am i doing that to the whole b because then ill end up with three columns instead of 2.

OpenStudy (unklerhaukus):

another example \[\mathrm M = \begin{bmatrix}-1&-2&-3\\2&3&0\end{bmatrix}\] \[-4\mathrm M = \begin{bmatrix}4&8&12\\-8&-12&0\end{bmatrix}\]

OpenStudy (unklerhaukus):

multiply every element of the matrix B, by -3

OpenStudy (anonymous):

Ok so i got -21 -15 -9 -12 6 -9

OpenStudy (unklerhaukus):

B = [7, 3, -2; 5, 4, 3] -3B = [(-3*7), (-3*3), (-3*-2); (-3*5), (-3*4), (-3*3)] = [-21, -9, 6; -15, -12, -9]

OpenStudy (unklerhaukus):

... Is there a typo in the question, or maybe the options?

OpenStudy (anonymous):

it was a typo in the answers the answer is a. thank you. @UnkleRhaukus

OpenStudy (unklerhaukus):

cool, d'you understand it all now?

OpenStudy (anonymous):

yeah after it was broken down. @UnkleRhaukus

OpenStudy (unklerhaukus):

Good \(\ddot\smile\)